NvidiaIntely (nvidiaintely):

Hard one coming! In the decomposition of Th-90, the mass defect = 4.50 x 10-24 kilograms. If all of this mass is converted to energy, what would be the value in joules? (1 J = 1 kg m2/s2) 4.05 x 10-7 J 1.50 x 10-8 J 1.35 x 10-16 J 5.4 x 10-32 J

NvidiaIntely (nvidiaintely):

@Kevin @Will.H @kittiwitti1

OpenStudy (will.h):

http://openstudy.com/updates/518d42e4e4b062a8d1da1715 They say it is A

NvidiaIntely (nvidiaintely):

I think so but, does @Kevin agree?

OpenStudy (kevin):

A

NvidiaIntely (nvidiaintely):

Okay, good again! ALL AGREES!

OpenStudy (kevin):

E = m x c^2 E = 4.50*10^-12 * (3 * 108 m/s2)^2

OpenStudy (will.h):

just use E=mc^2

NvidiaIntely (nvidiaintely):

Yeah